# Cohomogeneity one $\mathsf{RCD}$-spaces

### v2

@@ 2.5. Transformation groups
Let $(X,d,\mathfrak{m})$ be a metric measure space, and $G$ a topological group. A continuous (left) action of $G$ on $X$ is a continuous map $\alpha\colon G\times X\to X$ such that for any $g,h\in G$ and $x\in X$ we have $\alpha(g,\alpha(h,x))=\alpha(gh,x)$, and for $e\in G$ the identity element and any $x\in X$ we have $\alpha(e,x)=x$. The isotropy group at $x\in X$ is the closed subgroup $G_{x}=\left\{g\in G\mid gx=x\right\}$. We say the action $\alpha$ is effective if the only element of $G$ that fixes all the points in $X$ is the identity element, i.e. $\cap_{x\in X}G_{x}=\{e\}$. The action is proper if the map $A\colon G\times X\to X\times X$ given by $A(g,x)=(gx,x)$ is a proper map (that is, the preimage $A^{-1}(K)\subset G\times X$ of a compact subset $K\subset X\times X$ is compact. The orbit of a point $x\in X$ is the subset of $X$ given by $G(x)=\left\{gx\in X\mid g\in G\right\}$. The orbit type of an orbit $G(x)$ is the set $(G_{x}):=\{gG_{x}g^{-1}\mid g\in G\}$ of the conjugation class of the isotropy.
@@
When we have that $d(\alpha(g,x),\alpha(g,y))=d(x,y)$ for any $g\in G$ and any $x,y\in X$ we say that the action $\alpha$ is by isometries. Given $g\in G$ fixed we have a homeomorpshism $\alpha_{g}\colon X\to X$ given as $\alpha_{g}(x)=\alpha(g,x)$. In the case when the action is by isometries and $(\alpha_{g})_{\#}\mathfrak{m}=\mathfrak{m}$ for any $g\in G$, we say that the action is by measure-preserving isometries, and that the metric $d$ is $G$-invariant. For simplicity, unless otherwise stated, we denote the action as $gx:=\alpha(g,x)$. In the case the orbits are compact, the orbit space is equipped with the orbital metric
@@
The following theorem shows that a sufficient condition for the orbits to be compact is that the action is proper. Moreover, properness gives us the description of each orbit as a homogeneous space.
Theorem 2.26.
@@
In the case when $G$ is a compact Lie group, an action of $G$ on a metric space $X$ is proper (see for example [21, Theorem 2.2]). Thus from now on, we only consider compact Lie groups.
We recall from [27, Theorem 1.3] that given an $\mathsf{RCD}$-space $(X,d,\mathfrak{m})$ with an action by measure preserving isometries by a compact Lie group $G$, there exist (up to conjugation) a unique subgroup $G_{\min}<G$, such that the orbit $G(y)$ of $\mathfrak{m}$-a.e $y\in M$ is homeomorphic to the quotient $G/G_{\min}$. We call points $y\in M$ such that $G_{y}$ is conjugate to $G_{\min}$ in $G$ points with principal orbit type. The orbit of a point of principal orbit type is referred to as a principal orbit.
@@
We say $G$ acts Lipschitz and co-Lipschitz continuously on a metric space $(X,d)$, if for a bi-invariant metric on $G$, and every $y\in M$ with principal orbit type we have constants $C,R>0$ such that for all $0<r<R$ it holds
$B_{r/C}(y)\cap G(y)\subset\{gy\mid g\in B_{r}(e)\subset G\}\subset B_{Cr}(y)\cap G(y)$
In other words, the map $\alpha_{y}\colon G\to G(y)$ mapping $g$ to $gy$ is locally Lipschitz and co-Lipschitz continuous.
For an action of $G$ on $X$, given $x\in X$ fixed, a slice through $x$ is a subset $S_{x}\subset X$ that satisfies:
@@
    Setting $H=G_{x}$, there exists an open neighborhood $U\subset G/H$ around $eH$, and a cross-section $\chi\colon U\to G$ such that the function $F\colon U\times S_{x}\to X$ given by
    $F(gH,s)=\chi(gH)s,$
    is an homeomorphism onto its image.
We now state the main recognition tool in the theory of transformation groups, the so called Slice Theorem (see [10, Definition 4.1, Theorem 4.2]).
@@
Let a compact Lie group $G$ act by isometries on a metric space $(X,d)$. For all $x\in X$ for which a slice $S_{x}$ exists, there is some $r_{0}>0$ such that for all $r<r_{0}$ there is an equivariant homeomorphism
$G\times_{G_{x}}S_{x}\to B_{r}(G(x))$
Every point in a completely regular space has a slice (see [65]). This fact is also referred to sometimes as the Slice Theorem. It follows then that every point on an $\mathsf{RCD}$-space admits a slice. However, given an Alexandrov space $X$ and $G$ a Lie group acting by isometries, for each point $x\in X$ we can find slice containing $x$ which is a metric cone over a positively curved Alexandrov space (namely the space of normal directions at the point in question; see the the Slice Theorem for Alexandrov spaces due to Harvey and Searle, [44, Theorem B]). This implies that for any point in an Alexandrov space equipped with a compact Lie group action by isometries, there is a slice through the given point which in itself admits an Alexandrov structure.
Moreover, for the case when the space is a Riemannian manifold with a compact Lie group action by isometries the same holds: for any point there exists a slice containing the point, such that it admits a Riemannian space structure.
Thus in general, when considering a geometric space and a group action that preserve the geometric structure, we are interested in finding in a well defined way a slice which belongs to the same class of geometric objects as the global space.
@@ 2.6. Equivariant Gromov-Hausdorff convergence
Let $(X,d,x,G),$ $(Y,d_{Y},y,H)$ be two pointed metric spaces with $G$, $H$ groups acting isometrically on $X$ and $Y$ respectively. An equivariant $\epsilon$-Gromov-Hausdorff approximation is a triple $(f_{\epsilon},\varphi,\psi)$ which consists of maps
@@
    $d_{Y}(f_{\epsilon}(gp),\varphi(g)f_{\epsilon}(p)).$
@@ 3. Slice Theorem for cohomogeneity one geodesic spaces
In this Section we characterize the slices at each point of a compact, metric space $(X,d)$ admitting an action of a compact Lie group $G$ by isometries, having $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$ as orbit space, in terms of certain sets from the theory of optimal transport.
It is worth pointing out, however that we make no assumption of curvature conditions in this Section. To proceed, we recall a few notions from the theory of optimal transport (see for example the survey [16] for a brief account of the theory).
For each Lipschitz function $u\colon X\to\mathbb{R}$, we consider the subsets of $X\times X$ given by
$\Gamma_{u}$ :=$:=\left\{(x,y)\mid u(x)-u(y)=d(x,y)\right\},$
$\Gamma^{-}_{u}$ :=$:=\left\{(x,y)\mid(y,x)\in\Gamma_{u}\right\}.$
From these, we also define
$\mathcal{T}_{u}=\mathrm{pr}_{1}(\Gamma_{u}\smallsetminus\{x=y\})\cup\mathrm{pr}_{1}(\Gamma^{-}_{u}\smallsetminus\{x=y\}),$
where $\mathrm{pr}_{1}\colon X\times X\to X$ is given by $\mathrm{pr}_{1}(x,y)=x$. Now, for each $x\in X$ we set
$\Gamma_{u}(x)$ =$=\{y\in X\mid(x,y)\in\Gamma_{u}\},$
$\Gamma^{-}_{u}(x)$ =$=\{y\in X\mid(x,y)\in\Gamma^{-}_{u}\},$
and define an equivalence relation $R_{u}=\Gamma_{u}\cup\Gamma_{u}^{-}$.
We now consider the sets
$A_{+}(u)$ =$=\{y\in\mathcal{T}_{u}\mid\exists z,w\in\Gamma_{u}(y),\ (z,w)\not\in R_{u}\},$
$A_{-}(u)$ =$=\{y\in\mathcal{T}_{u}\mid\exists z,w\in\Gamma_{u}^{-}(y),\ (z,w)\not\in R_{u}\}.$
The non-branching transport set of $u$ is defined as $\mathcal{T}^{nb}_{u}=\mathcal{T}_{u}\setminus(A_{+}(u)\cup A_{-}(u))$.
We now state the main result of this Section, containing the definition of our distinguished slices.
Theorem C.
@@
Let $(X,d)$ be a geodesic space and $G$ a compact Lie group acting on $X$ by isometries such that $X^{\ast}$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$. Then for any $x_{0}\in X$ we define $u(y):=d(y,G(x_{0}))$, and there exists $\delta_{0}>0$ such that for any $0<\delta<\delta_{0}$ we have $B_{\delta}(x_{0})\subset\mathcal{T}^{nb}_{u}$ and
$S_{x_{0}}:=\left(\Gamma^{-}_{u}(x_{0})\cap B_{\delta}(x_{0})\right)\cup\{x_{0}\}$
@@
We point out that $S_{x_{0}}$ is homeomorphic to a topological cone, since it consists of a union of geodesics, such that the only common point of two geodesics is the vertex.
To show the validity of Theorem C, we need to prove some results first. Assume now that we have a compact Lie group $G$ acting effectively by isometries on $(X,d)$. Let $u\colon X\to\mathbb{R}$ be a Lipschitz function which is $G$-invariant, i.e. for $x\in X$ we have $u(gx)=u(x)$ for any $g\in G$. Observe that we have a well defined Lipschitz function $u^{\ast}\colon(X/G,d^{\ast})\to\mathbb{R}$ given by $u^{\ast}(x^{\ast}):=u(x)$. Observe that $u^{\ast}$ is well defined, since for $g\in G$ we have that $u(gx)=u(x)$ by hypothesis. Thus $u^{\ast}(x^{\ast})$ does not depend on the representative $x$ of $x^{\ast}$.
@@
Let $x_{0}\in X$ be fixed and $u\colon X\to\mathbb{R}$ be the $G$-invariant Lipschitz function given by $u(y)=d(G(x_{0}),y)$. Then
$\left\{(x^{\ast},y^{\ast})\mid(x,y)\in\Gamma_{u}\right\}\subseteq\Gamma_{u^{\ast}}.$
Moreover, for every $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$, there exists $g\in G$ such that $(gx,y)\in\Gamma_{u}$ and $d^{\ast}(x^{\ast},y^{\ast})=d(gx,y)$. In an analogous fashion, due to the symmetry in the definition of $\Gamma_{u}^{-}$, the same holds for $\Gamma_{u}^{-}$.
@@
Consider $(x,y)\in\Gamma_{u}$. Observe that $d(G(x_{0}),x)=d^{\ast}(x_{0}^{\ast},x^{\ast})$, and thus
$d^{\ast}(x^{\ast},y^{\ast})\leq d(x,y)$ $=u(x)-u(y)$
$=u^{\ast}(x^{\ast})-u^{\ast}(y^{\ast})$
$=d^{\ast}(x_{0}^{\ast},x^{\ast})-d^{\ast}(x_{0}^{\ast},y^{\ast})$
$\leq d^{\ast}(x_{0}^{\ast},y^{\ast})+d^{\ast}(y^{\ast},x^{\ast})-d^{\ast}(x_{0}^{\ast},y^{\ast})$
$=d^{\ast}(x^{\ast},y^{\ast}).$
Thus we conclude that $d^{\ast}(x^{\ast},y^{\ast})=u^{\ast}(x^{\ast})-u^{\ast}(y^{\ast})$, and therefore $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$, proving the first claim.
We now consider $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$. Then we have that for any $\bar{g},\bar{h}\in G$ it holds that
$d^{\ast}(x^{\ast},y^{\ast})=u^{\ast}(x^{\ast})-u^{\ast}(y^{\ast})=u(\bar{g}x)-u(\bar{h}y).$
Also observe that there exists $g\in G$ such that $d^{\ast}(x^{\ast},y^{\ast})=d(gx,y)$. Thus by setting $\bar{g}=g$ and $\bar{h}=e$ we have
$d(gx,y)=d^{\ast}(x^{\ast},y^{\ast})=u(gx)-u(y).$
This implies that $(gx,y)\in\Gamma_{u}$. ∎

### v3

@@ 2.5. Transformation groups
Let $(X,d,\mathfrak{m})$ be a metric measure space, and $G$ a topological group. A continuous (left) action of $G$ on $X$ is a continuous map $\alpha\colon G\times X\to X$ such that for any $g,h\in G$ and $x\in X$ we have $\alpha(g,\alpha(h,x))=\alpha(gh,x)$, and for $e\in G$ the identity element and any $x\in X$ we have $\alpha(e,x)=x$. From now on we denote $gx:=\alpha(g,x)$.
The isotropy group at $x\in X$ is the closed subgroup $G_{x}=\left\{g\in G\mid gx=x\right\}$. We say the action $\alpha$ is effective if the only element of $G$ that fixes all the points in $X$ is the identity element, i.e. $\cap_{x\in X}G_{x}=\{e\}$. The action is proper if the map $A\colon G\times X\to X\times X$ given by $A(g,x)=(gx,x)$ is a proper map (that is, the preimage $A^{-1}(K)\subset G\times X$ of a compact subset $K\subset X\times X$ is compact.
The orbit of a point $x\in X$ is the subset of $X$ given by $G(x)=\left\{gx\in X\mid g\in G\right\}$. The orbit type of an orbit $G(x)$ is the conjugation class $(G_{x}):=\{gG_{x}g^{-1}\mid g\in G\}$ of the isotropy.
@@
When we have that $d(\alpha(g,x),\alpha(g,y))=d(x,y)$ for any $g\in G$ and any $x,y\in X$ we say that the action $\alpha$ is by isometries. Given $g\in G$ fixed we have a homeomorphism $\alpha_{g}\colon X\to X$ given as $\alpha_{g}(x)=\alpha(g,x)$. In the case when the action is by isometries and $(\alpha_{g})_{\#}\mathfrak{m}=\mathfrak{m}$ for any $g\in G$, we say that the action is by measure-preserving isometries, and that the metric $d$ is $G$-invariant. For simplicity, unless otherwise stated, we denote the action as $gx:=\alpha(g,x)$. In the case where the orbits are compact, the orbit space is equipped with the orbital metric
@@
When the action is proper, the isotropy groups are compact. Moreover, properness gives us the description of each orbit as a homogeneous space.
Theorem 2.24 (Theorem 2.3 in [20]).
@@
In the case when $G$ is a compact Lie group, an action of $G$ on a metric space $X$ is proper (see for example [20, Theorem 2.2]). From now on, we only consider compact Lie groups.
We recall from [25, Theorem 1.3] that given an $\mathsf{RCD}$-space $(X,d,\mathfrak{m})$ with an action by measure preserving isometries by a compact Lie group $G$, there exist (up to conjugation) a unique subgroup $G_{\min}<G$, such that the orbit $G(y)$ of $\mathfrak{m}$-a.e $y\in M$ is homeomorphic to the quotient $G/G_{\min}$. We call points $y\in M$ such that $G_{y}$ is conjugate to $G_{\min}$ in $G$ points with principal orbit type. The orbit of a point of principal orbit type is referred to as a principal orbit.
@@
We will use the following definition of slice taken from Bredon [11]
@@
Let $x\in X$ a set $x\in S$ is called a slice at $x$ if it satisfies the following:
@@
    $G_{x}(S)=S$
  4. (iv)
    $(gS)\cap S\neq\emptyset$ implies that $g\in G_{x}$.
Sometimes we will use the following equivalent characterization of a slice.
Theorem 2.27 (Chapter $2$, Section $4$ in [11]).
For an action $G$ on $X$, given a point $x\in X$ a slice through $x$ can be equivalently defined as a subset $S_{x}\subset X$ that satisfies:
  • $gS_{x}\subset S_{x}$ for all $g\in G_{x}$
  • $gS_{x}\cap S_{x}\neq\emptyset$ implies that $g\in G_{x}$
  • Setting $H=G_{x}$, there exists an open neighborhood $U\subset G/H$ around $eH$, and a cross-section $\chi\colon U\rightarrow G$ such that the function $F\colon U\times S_{x}\rightarrow X$ given by
    $F(gH,s)=\chi(gH)s$
    is a homeomorphism onto its image.
We now state the main recognition tool in the theory of transformation groups, the so called Slice Theorem (see [11, Definition 4.1, Theorem 4.2]).
@@
Let a compact Lie group $G$ act by isometries on a metric space $(X,d)$. For all $x\in X$ there is some $r_{0}>0$, such that for all $0<r\leq r_{0}$ there exists a slice $S_{x}=S_{x}(r)$ and there is an equivariant homeomorphism
$G\times_{G_{x}}S_{x}\to B_{r}(G(x)).$
Remark 2.29.
The dependence of $S_{x_{0}}$ on the radius $r$ is given by ?THM? LABEL:MT:_Slice_Theorem_Cohomogenity_one.
Every point in a completely regular space has a slice (see [68]). This fact is sometimes referred to as the Slice Theorem. It follows then that every point on an $\mathsf{RCD}$-space admits a slice. However, given an Alexandrov space $X$ and $G$ a Lie group acting by isometries, for each point $x\in X$ we can find a slice containing $x$ which is a metric cone over a positively curved Alexandrov space (namely the space of normal directions at the point in question; see the Slice Theorem for Alexandrov spaces due to Harvey and Searle, [42, Theorem B]). This implies that for any point in an Alexandrov space equipped with a compact Lie group action by isometries, there is a slice through the given point which in itself admits an Alexandrov structure.
Moreover, for the case when the space is a Riemannian manifold with a compact Lie group action by isometries, the same holds: for any point there exists a slice containing the point, such that it admits a Riemannian space structure.
Thus in general, when considering a geometric space and a group action that preserves the geometric structure, we are interested in finding in a well-defined way a slice which belongs to the same class of geometric objects as the global space.
@@ 2.6. Equivariant Gromov-Hausdorff convergence
Let $(X,d_{X},x,G)$, $(Y,d_{Y},y,H)$ be two pointed metric spaces with $G$, $H$ groups acting isometrically on $X$ and $Y$ respectively. An equivariant $\epsilon$-Gromov-Hausdorff approximation is a triple $(f_{\epsilon},\varphi,\psi)$ which consists of maps
@@
    $d_{Y}(f_{\epsilon}(gp),\varphi(g)f_{\epsilon}(p))<\epsilon.$
@@
With this definition we present the following result that is used later on.
Theorem 2.32 (Proposition 3.6 in [24]).
Assume that for a sequence $\{X_{i},d_{X_{i}},H_{i},x_{i})\}_{i\in\mathbb{N}}$ of pointed metric spaces with $H_{i}\subset\mathrm{Isom}(X_{i},d_{i})$ a closed subgroup with respect to the compact-open topology acting effectively on $X_{i}$, we we have that $(Y,d_{Y},y)$ is the pointed-Gromov-Hausdorff limit of $\{(X_{i},d_{X_{i}},x_{i})\}_{i\in\mathbb{N}}$. Then there exist $H^{\prime}\subset\mathrm{Isom}(Y,d_{Y})$ a closed subgroup with respect to the compact-open topology, acting effectively on $Y$, and a subsequence $\{(X_{i_{k}},d_{X_{i_{k}}},H_{i_{k}},x_{i_{k}})\}_{k\in\mathbb{N}}$ converging to $(Y,d_{Y},H^{\prime},y)$ in the pointed-equivariant Gromov-Hausdorff sense.
@@ 3. Slice Theorem for cohomogeneity one geodesic spaces
In this section we do not need any curvature bounds on $X$. However, we need to assume some properties for some geodesics of $X$. In order to properly describe these properties we need to introduce some concepts of optimal transport which we recall below. The interested reader can find the proofs of the claims for example in [4].
A metric space $(X,d)$ is called a geodesic space if any two points $x_{0},x_{1}\in X$ can be joined by a curve $\gamma\colon[0,1]\rightarrow X$ with $\gamma_{0}=x_{0},\gamma_{1}=x_{1}$ and such that:
$d(\gamma_{t},\gamma_{s})=|t-s|d(\gamma_{0},\gamma_{1}),\text{ for all }s,t\in[0,1].$
The space of all geodesics in $X$ is denoted by $\operatorname{Geo}(X)$ and equipped with the $\sup$ norm.
Let $\mathbb{P}(X)$ denote the space of Borel probability measures on $X$, and for $p\in[1,\infty)$ by $\mathbb{P}_{p}(X)$ those that have finite $p$-moments, that is
$\mathbb{P}_{p}(X):=\Big\{\mu\in\mathbb{P}(X)\,|\,\int d^{p}(x,x_{0})d\mu(x)<\infty\text{ for some }x_{0}\in X\Big\}.$
We then endow $\mathbb{P}_{p}(X)$ with the $L^{p}$-Wasserstein distance: For all $\mu_{0},\mu_{1}\in\mathbb{P}_{p}(X)$
(4) $\mathbb{W}_{p}^{p}(\mu_{0},\mu_{1})=\inf_{\rho}\int d^{p}(x,y)d\rho(x,y),$
where the infimum is taken over all probability measures in $\rho\in\mathbb{P}(X\times X)$ such that their marginals are $\mu_{0}$ and $\mu_{1}$, i.e. for the projections $\times X\ni(x,y)\mapsto\mathrm{proj}_{1}(x,y)=x, X×X$$\times X\ni(x,y)\mapsto\mathrm{proj}_{2}(x,y)=y$ we have $(\mathrm{proj}_{i})_{\#}\rho=\mu_{i-1}$ for $i\in\{1,2\}$. It can be proved that one can always find measures that achieve such infimum, they are called optimal transports and we denote the set of them by $\operatorname{Opt}(\mu_{0},\mu_{1})$.
The $L^{2}$-Wasserstein space will inherit some geometric properties from its base. For example, if $(X,d)$ is a geodesic space, then $(\mathbb{P}_{2}(X),\mathbb{W}_{2})$ is also a geodesic space. Furthermore, the next result tell us that the optimal transport between two measures occurs by moving along geodesics of $X$. Recall that for $t\in[0,1]$, the evaluation map $e_{t}\colon\operatorname{Geo}(X)\rightarrow X$ is defined as $e_{t}(\gamma)=\gamma_{t}$.
Theorem 3.1.
(Theorem 3.10 in [4]) Let $(X,d)$ be a Polish geodesic space. Then the $L^{2}$-Wasserstein space $(\mathbb{P}_{2}(X),\mathbb{W}_{2})$ is also a geodesic space. Furthermore, the following are equivalent:
  1. (1)
    A curve $t\mapsto\eta_{t}$ is a geodesic.
  2. (2)
    There exists a measure $\eta\in\mathbb{P}(\operatorname{Geo}(X))$ called a dynamical optimal transport, such that $(e_{0},e_{1})_{\#}\eta\in\operatorname{Opt}(\eta_{0},\eta_{1})$ and
    $\eta_{t}=e_{t\#}\eta.$
The set of measures described in $(2)$ is denoted by $\operatorname{Opt}\operatorname{Geo}(\mu_{0},\mu_{1})$.
We recall the definition of an essentially non-branching space which was first introduced by Rajala and Sturm in [72]. Although in said paper the authors were working with spaces satisfying a lower Ricci curvature bound, we stress that the definition does not rely on it.
Definition 3.2.
A metric measure space $(X,d,\mathfrak{m})$ is essentially non-branching if for every $\mu_{0},\mu_{1}\in\mathbb{P}_{2}(X)$ such that $\mu_{0},\mu_{1}\ll\mathfrak{m}$, we have that every $\rho\in\operatorname{Opt}\operatorname{Geo}(\mu_{0},\mu_{1})$ is concentrated on a set of non-branching geodesics.
Remark 3.3.
In [67] Ohta constructed a couple of explicit examples of spaces that are essentially non-branching in the sense of ?THM? LABEL:D:_essentiallynonbranch and that do exhibit some branching geodesics.
We now recall the main theorem of this section:
Theorem LABEL:MT:_Slice_Theorem_Cohomogenity_one.
@@
Let $(X,d,\mathfrak{m})$ be an essentially non-branching space and $G$ a compact Lie group acting on $X$ by isometries such that $X^{\ast}$ is isometric to [1,1][-1,1], $0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$. Then for any $x_{0}\in X$, there exists $\delta_{0}>0$ such that, for any $0<\delta<\delta_{0}$ we have that
$S_{x_{0}}:=\{y\in X\,|\,d(x_{0},y)=d(G(x_{0}),y)\}\cap B_{\delta}(x_{0})$
@@
Here we need to take $\delta_{0}$ small enough so that $\pi(B_{\delta_{0}}(x_{0}))$ is homeomorphic to one of $(x_{0}^{\ast}-\delta_{0},x_{0}^{\ast}+\delta_{0})$ or $x_{0}^{\ast},x_{0}^{\ast}+\delta)$ or $(x_{0}-\delta_{0},x_{0}]$, for $\pi\colon X\to X/G$.
Before going through the proof of this Theorem we will need to introduce a particular class of geodesics which let us define the set that turns to be the slice.
Definition 3.5.
Let $\pi\colon X\rightarrow X^{\ast}$ be the quotient map. We will say that a geodesic $\gamma\in\operatorname{Geo}(X)$ is horizontal if $\pi\circ\gamma\in\operatorname{Geo}(X^{\ast})$.
@@
Let $\gamma\in\operatorname{Geo}(X)$ be such that $d(\gamma_{0},\gamma_{1})=d^{\ast}(\gamma_{0}^{\ast},\gamma_{1}^{\ast})$. Then $\pi\circ\gamma\in\operatorname{Geo}(X^{\ast})$.
@@
Let $\gamma\in\operatorname{Geo}(X)$ be such that $d(\gamma_{0},\gamma_{1})=d^{\ast}(\gamma_{0}^{\ast},\gamma_{1}^{\ast})$, and denote by $\gamma^{\ast}=\pi\circ\gamma$. Take $s,t\in[0,1]$, and observe that we can assume without loss of generality that $s<t$. Now there exists some $g\in G$ such that $d(\gamma_{s},g\gamma_{t})=d^{\ast}(\gamma_{s}^{\ast},\gamma_{t}^{\ast})$. It is clear that $d(g\gamma_{t},g\gamma_{1})=d(\gamma_{t},\gamma_{1})$ so then
$d^{\ast}(\gamma_{0}^{\ast},\gamma_{1}^{\ast})$ $\leq d(\gamma_{0},\gamma_{1})$
$\leq d(\gamma_{0},\gamma_{s})+d(\gamma_{s},g\gamma_{t})+d(g\gamma_{t},g\gamma_{1})$
$\leq d(\gamma_{0},\gamma_{s})+d(\gamma_{s},\gamma_{t})+d(\gamma_{t},\gamma_{1})$
$=d(\gamma_{0},\gamma_{1})=d^{\ast}(\gamma_{0}^{\ast},\gamma_{1}^{\ast}).$
This implies that $d(\gamma_{s},\gamma_{t})=d(\gamma_{s},g\gamma_{t})=d^{\ast}(\gamma_{s}^{\ast},\gamma_{t}^{\ast})$. ∎